When UV light of wavelength 243nm falls on a metal surface, the maximum kinetic energy of emitted electrons is 1.46eV. What is the work function of the metal?
Answer #1
hc/wavelength=work function+KE
6.67*10-34*3*10^8/243*10-9=W+1.46*1.6*10-19
W=5.898*10-19 J